# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ 5.2 Algorithm, Theorem and Proof Sketch
Our algorithm, Contextual OLBACID (Algorithm 3), or COLBACID in short, works as follows. In period $t$, we first compute a ridge estimator $\hat{\boldsymbol{\theta}}(t)$ of $\boldsymbol{\theta}^{\star}$ by (13). The algorithm then classifies a new type-$k$ post as harmful ($Y_{k}(t)=0$) if its empirical average cost, $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}(t)$, is positive. We follow the same label-driven admission with OLBACID in Line 3 but we set the confidence interval on $h_{k}$ by (15) using the confidence set $\mathcal{C}_{t}$ from (14). The optimistic admission rule (Line 3) similarly finds an optimistic per-period idiosyncrasy loss $\bar{r}_{k}(t)$ but estimates the delay loss by type-aggregated queue lengths. Specifically, letting $\tilde{\mathcal{Q}}_{g}(t)=\{j\in\mathcal{Q}(t)\colon g(k(j))=g\}$ be the set of waiting posts whose types belong to group $g$ and $\tilde{Q}_{g}(t)=|\tilde{\mathcal{Q}}_{g}(t)|$, we admit a type-$k$ post if and only if its (optimistic) idiosyncrasy loss is higher than the estimated delay loss, i.e., $\beta\bar{r}_{k(t)}(t)\ell_{k}\geq\tilde{Q}_{g(k(t))}(t)$. We still prioritize the label-driven queue for scheduling. If there is no post in the label-driven queue, we use a type-aggregated MaxWeight scheduling: we first pick a group $g$ maximizing $\tilde{\mu}_{g}\tilde{Q}_{g}(t)$, and then pick the earliest admitted post in the review queue $\mathcal{Q}(t)$ whose type is of group $g$ to review.
Data: $(\ell_{k},\mu_{k},\sigma_{k}\}_{k\in\mathcal{K}}$, upper bounds $U,H$, and group partition $\mathcal{K}_{\mathcal{G}}$
1 $\beta,\gamma\leftarrow 1/\sqrt{G\ell_{\max}},~{}\delta\leftarrow\min(\gamma,1/T),~{}\kappa\leftarrow\max(1,U^{2})$
@@ 3       Observe a new post $j(t)=j$ of type $k(j)=k$; if no new post, set $j(t)=\perp$
4       Compute ridge estimator $\hat{\boldsymbol{\theta}}(t)$ by (13)
5       if $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}(t)>0$ then $Y_{k}(t)\leftarrow 0$ else $Y_{k}(t)\leftarrow 1$ // Empirical Classification
@@ 4       Compute ridge estimator $\hat{\boldsymbol{\theta}}(t)$ by (13)
6       $\bar{r}_{k}(t)=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)$ where $\underline{h}_{k}(t),\bar{h}_{k}(t)$ are given by (15)
@@
For ease of exposition, we follow the $\lesssim$ notation to include only super-logarithmic dependence on the number of groups $G$, feature dimension $d$, maximum lifetime $\ell_{\max}$, the margin $\eta$, window size $w$, the time horizon $T$, and the aggregation gap $\Delta(\mathcal{K}_{\mathcal{G}})$. Our main result is as follows.
@@
Similar to the proof of Theorem 3, we set $v_{\max}=\max_{k}v_{k},\sigma_{\max}=\max(1,\max_{k}\sigma_{k})$, $\hat{\mu}_{\min}=\min_{t\leq T,k\in\mathcal{K}}N(t)\mu_{k}$, $r_{\max}=v_{\max}(H+\sigma_{\max})\geq 1$ and focus on $T\geq 3$ and $U,H\geq 1$. The following lemmas also assume $G\leq\ell_{\max}$; Theorem 4 holds directly otherwise.
@@
For any $\beta,\gamma\in(1/\ell_{\max},1),\kappa\geq U^{2}$ and $\delta\in(0,1/T)$,
$\text{Relaxed Delay Loss}\leq 2G\beta r^{2}_{\max}\ell_{\max}T+\frac{9r_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$
@@
For any $\beta,\gamma\in(1/\ell_{\max},1]$, $\kappa\geq U^{2}$ and $\delta\in(0,\min(1/T,\gamma))$,
$\text{Classification Loss}\leq 4\gamma r_{\max}\ell_{\max}T+\frac{9r_{\max}\ell_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$
@@
For any $\beta,\gamma\in(1/\ell_{\max},1]$, $\kappa\geq U^{2}$ and $\delta\in(0,1/T)$,
@@
Similar to Section 4.3, we bound the relaxed delay loss by the sum of the label-driven queue length $\mathbb{E}\left[\sum_{t=1}^{T}Q^{\mathcal{\textsc{LD}}}(t)\right]$ and the review queue length $\mathbb{E}\left[\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}\tilde{Q}_{g}(t)\right]$. By Lemma 4.4, bounding the first term requires bounding the the expected number of posts that we admit into the label-driven queue, i.e., $\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\right].$ We first define the “good” event $\mathcal{E}$, where the confidence set $\mathcal{C}_{t}$ is valid for any period $t$: $\mathcal{E}\coloneqq\{\forall t,\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t}\}$. Our first lemma shows that this event happens with probability at least $1-\delta$ using [APS11, Theorem 2] (proof in Appendix D.3).
@@
For any $\delta\in(0,1)$, with probability at least $1-\delta$, for any $t\geq 0$, the confidence set $\mathcal{C}_{t}$ is valid, i.e., $\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t}$.
@@
If $\kappa\geq U^{2},\gamma\leq 1$, $\delta\leq 1/T$, the number of posts admitted to the label-driven queue is
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\right]\leq\frac{9B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}.$
@@
$\leq r_{\max}\left(2GT\beta r_{\max}\ell_{\max}+\frac{9B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}\right)=2GT\beta r^{2}_{\max}\ell_{\max}+\frac{9r_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$
@@
For any type $k$ and period $t$, $Z_{k}(t)\mathbbm{1}\left(\mathcal{E}\right)\leq\left(3v_{\max}\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\Lambda_{k}(t).$
@@
$\leq r_{\max}\ell_{\max}\delta T+\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}Z_{k}(t)\,\middle|\,\mathcal{E}\right]$ (By Lemma 5.4)
$\leq r_{\max}\ell_{\max}\gamma T+\mathbb{E}\left[\sum_{t=1}^{T}\left(3v_{\max}\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\right]$ (By Lemma 5.7 and $\delta\leq\gamma$)
$\leq 4\gamma r_{\max}\ell_{\max}T+\frac{9r_{\max}\ell_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$ (By Lemmas 4.45.5)
@@
If $\kappa\geq U^{2},\gamma\leq 1,\delta\leq 1/T$, the regret in scheduling is $\textsc{RegS}(T)\leq\frac{9Q_{\max}B^{2}_{\delta}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$
@@
If $\kappa\geq U^{2},\gamma\leq 1,\delta\leq 1/T$, the regret in admission is
$\textsc{RegA}(T)\leq 3\beta r_{\max}\ell_{\max}B_{\delta}(T)\left(d\ln(1+T/d)\left(4GQ_{\max}+\frac{9B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right)+\sqrt{2GTd\ln(1+T/d)}\right).$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}\bar{A}_{k}(t)\right]\leq\sqrt{2GTd\ln(1+T/d)}+d\ln(1+T/d)\left(3GQ_{\max}+\frac{9B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right).$

### v3

@@ 5.2 Algorithm, Theorem and Proof Sketch
Our algorithm, Contextual OLBACID (Algorithm 3), or COLBACID in short, works as follows. In period $t$, we first compute a ridge estimator $\hat{\boldsymbol{\Theta}}(t-1)$ of $\boldsymbol{\Theta}^{\star}$ by (14). The algorithm then classifies a new type-$k$ post as harmful ($Y_{k}(t)=0$) if its empirical average cost, $\hat{h}_{k}(t)=\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{h}(t-1)$, is positive. We follow the same label-driven admission with OLBACID in Line 3 but we set the confidence interval on $h_{k}$ by (16) using the confidence set $\mathcal{C}_{t}$ from (15). The optimistic admission rule (Line 3) similarly finds an optimistic per-period idiosyncrasy loss $\bar{r}_{k}(t)$ but estimates the delay loss by type-aggregated queue lengths. Specifically, letting $\tilde{\mathcal{Q}}_{g}(t)=\{j\in\mathcal{Q}(t)\colon g(k(j))=g\}$ be the set of waiting posts whose types belong to group $g$ and $\tilde{Q}_{g}(t)=|\tilde{\mathcal{Q}}_{g}(t)|$, we admit a type-$k$ post if and only if its (optimistic) idiosyncrasy loss is higher than the estimated delay loss, i.e., $\beta\bar{r}_{k(t)}(t)\ell_{k}\geq\tilde{Q}_{g(k(t))}(t)$. We still prioritize the label-driven queue for scheduling. If there is no post in the label-driven queue, we use a type-aggregated MaxWeight scheduling: we first pick a group $g$ maximizing $\tilde{\mu}_{g}\tilde{Q}_{g}(t)$, and then pick the earliest admitted post in the review queue $\mathcal{Q}(t)$ whose type is of group $g$ to review.
Data: $(\ell_{k},\mu_{k}\}_{k\in\mathcal{K}}$, upper bounds $U,r_{\max},\sigma_{\max}$, and group partition $\mathcal{K}_{\mathcal{G}}$
Admission parameters: $\beta,\gamma\leftarrow 1/\sqrt{G\ell_{\max}},~{}\delta\leftarrow\min(\gamma,0.5/T),~{}\kappa\leftarrow\max(1,U^{2})$
@@ 2       Observe a new post $j(t)=j$ of type $k(j)=k$; if no new post, set $j(t)=\perp$
3       Compute ridge estimator $\hat{\boldsymbol{\Theta}}(t-1)$ by (14)
4       if $\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{h}(t-1)>0$ then $Y_{k}(t)\leftarrow 0$ else $Y_{k}(t)\leftarrow 1$ // Empirical Classification
@@ 3       Compute ridge estimator $\hat{\boldsymbol{\Theta}}(t-1)$ by (14)
5       Calculate $\underline{h}_{k}(t),\bar{h}_{k}(t),\bar{r}_{k}(t)$ by (16) and (17)
@@
For ease of exposition, we follow the $\lesssim$ notation to include only super-logarithmic dependence on the number of groups $G$, feature dimension $d$, maximum lifetime $\ell_{\max}$, the margin $\eta$ ($\eta=\min(1,\min_{k\in\mathcal{K}}|h_{k}|)$, window size $w$, the time horizon $T$, and the aggregation gap $\Delta(\mathcal{K}_{\mathcal{G}})$. Our main result is as follows.
@@
Similar to the proof of Theorem 3, we set $\hat{\mu}_{\min}=\min_{t\leq T,k\in\mathcal{K}}N(t)\mu_{k}$ and focus on $T\geq 3$. The following lemmas also assume $G\leq\ell_{\max}$; Theorem 4 holds directly otherwise.
@@
For any $\beta,\gamma\in(1/\ell_{\max},1),\kappa\geq U^{2}$ and $\delta\in(0,0.5/T)$,
$\text{Relaxed Delay Loss}\leq 2G\beta r^{2}_{\max}\ell_{\max}T+\frac{34r_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$
@@
For any $\beta,\gamma\in(1/\ell_{\max},1]$, $\kappa\geq U^{2}$ and $\delta\in(0,\min(0.5/T,\gamma))$,
$\text{Classification Loss}\leq 3\gamma r_{\max}\ell_{\max}T+\frac{34r_{\max}\ell_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$
@@
For any $\beta,\gamma\in(1/\ell_{\max},1]$, $\kappa\geq U^{2}$ and $\delta\in(0,0.5/T)$,
@@
Similar to Section 4.3, we bound the relaxed delay loss by the sum of the label-driven queue length $\mathbb{E}\left[\sum_{t=1}^{T}Q^{\mathcal{\textsc{LD}}}(t)\right]$ and the review queue length $\mathbb{E}\left[\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}\tilde{Q}_{g}(t)\right]$. By Lemma 4.4, bounding the first term requires bounding the the expected number of posts that we admit into the label-driven queue, i.e., $\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\right].$ We first define the “good” event $\mathcal{E}$, where the confidence set $\mathcal{C}_{t}$ is valid for any period $t$: $\mathcal{E}\coloneqq\{\forall t,\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t}\}$. Our first lemma shows that this event happens with probability at least $1-2\delta$ using [APS11, Theorem 2] (proof in Appendix D.3).
@@
For any $\delta\in(0,1)$, it holds that with probability at least $1-2\delta$, for any $t\geq 0$, the confidence set $\mathcal{C}_{t}$ is valid, i.e., $\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t}$.
@@
If $\kappa\geq U^{2},\gamma\leq 1$, $\delta\leq 0.5/T$, the number of posts admitted to the label-driven queue is
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\right]\leq\frac{34B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}.$
@@
$\leq r_{\max}\left(2GT\beta r_{\max}\ell_{\max}+\frac{34B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}\right)=2GT\beta r^{2}_{\max}\ell_{\max}+\frac{34r_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$
@@
For any type $k$ and period $t$, $Z_{k}(t)\mathbbm{1}\left(\mathcal{E}\right)\leq\left(\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\Lambda_{k}(t).$
@@
$\leq 2r_{\max}\ell_{\max}\delta T+\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}Z_{k}(t)\,\middle|\,\mathcal{E}\right]$ (By Lemma 5.4)
$\leq 2r_{\max}\ell_{\max}\gamma T+\mathbb{E}\left[\sum_{t=1}^{T}\left(\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\right]$ (By Lemma 5.7 and $\delta\leq\gamma$)
$\leq 3\gamma r_{\max}\ell_{\max}T+\frac{34r_{\max}\ell_{\max}B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$ (By Lemmas 4.45.5)
@@
If $\kappa\geq U^{2},\gamma\leq 1,\delta\leq 0.5/T$, the regret in scheduling is $\textsc{RegS}(T)\leq\frac{68Q_{\max}B^{2}_{\delta}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}.$
@@
If $\kappa\geq U^{2},\gamma\leq 1,\delta\leq 0.5/T$, the regret in admission is
$\textsc{RegA}(T)\leq 3\beta r_{\max}\ell_{\max}B_{\delta}(T)\left(d\ln(1+T/d)\left(4GQ_{\max}+\frac{34B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right)+\sqrt{2GTd\ln(1+T/d)}\right).$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k}\|\boldsymbol{\phi}_{k}\|_{\bar{\boldsymbol{V}}_{t-1}^{-1}}\bar{A}_{k}(t)\right]\leq\sqrt{2GTd\ln(1+T/d)}+d\ln(1+T/d)\left(3GQ_{\max}+\frac{34B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right).$
@@
6 Numerical Results
We now provide numerical support to the efficiency of BACID compared with congestion-unaware heuristics, and to the need of label-driven admission to ensure correct classification decisions.
6.1 The Effectiveness of BACID in a Non-stationary Environment
We consider a non-stationary setting with $T=500000$ periods. There are two types of posts, with arrival rates $\lambda_{1}(t)\equiv 0.2,\lambda_{2}(t)\equiv 0.4$ and service rates $\mu_{1}=\mu_{2}=0.05$. The human capacity is time-varying in the sense that for every $5000$ periods, the capacity is high for the initial $4000$ periods with $N(t)=9$ while the capacity is low for the remaining $1000$ periods with $N(t)=2$. The cost distribution of type $k\in\{1,2\}$, $\mathcal{F}_{k}$, is a normal distribution with mean $h_{k}$ and standard deviation $1$ where $h_{1}=-1$ and $h_{2}=0.1$. Both types have equal lifetime $\ell$ that we specify later.
Assuming knowledge of $\mathcal{F}_{k}$, we next compare the efficiency of BACID with five congestion-unaware heuristics: AI-Only and Human-Only as given in Section 3.1, Static-$k$ with $k\in\{1,2\}$ that statically admits either type-$1$ posts or type-$2$ posts, and Dynamic, which, given the optimal solution $\boldsymbol{a}^{\star}(t)$ of the $1-$fluid benchmark, admits a type-$k$ post with probability $a^{\star}_{k}(t)/\lambda_{k}(t)$ in period $t$. Dynamic is thus a dynamic static policy that adjusts the set of types to admit in each period based on arrival rates and human capacities, but not based on real-time congestion information.
There are two sets of results as shown in Figures 4 and 4. In particular, for Figure 4, we fix the lifetime $\ell$ to be $500$ and plot, for each policy, its average regret for a window size $1$, $\textsc{Reg}^{\pi}(1,t)$ defined in (3), for $t\leq T$. Figure 4 shows that while the regret of BACID is consistently close to 00, the regrets of the congestion-unaware policies are much higher. To further illustrate the suboptimality of congestion-unaware policies, Figure 4 shows the average regret of all policies, $\textsc{Reg}^{\pi}(1,T)$, but as a function of the lifetime $\ell$, which varies from $50$ to $500$. Consistent with Section 3, Figure 4 shows that congestion-unaware policies in general have average regret linear in $\ell$, but the average regret of BACID is sublinear in $\ell$. Indeed, the average regret of BACID is close to 00 across all lifetimes and behaves better than the $O(\sqrt{\ell})$ bound established in Theorem 1.
Figure 3: $\textsc{Reg}^{\pi}(1,t)$ as a function of periods
Figure 4: $\textsc{Reg}^{\pi}(1,T)$ as a function of lifetime $\ell$
6.2 Comparison of OLBACID and Optimism-Only Learning
We next numerically explore scenarios where an optimism-only approach fails to learn classification decisions and therefore incurs large regret. Similar to the setting in Section 4.1, there are $T=10^{5}$ periods with $K=2$ types of posts. For a type-$1$ post, its cost is $-1$ or $+1$ with equal probability $0.5$. For a type-$2$ post, its cost is $-1$ with probability $0.05$ and $+1$ with probability $0.95$. The lifetimes are $\ell_{1}=10000$ and $\ell_{2}=1000$. Service rates are both $0.5$ and $N(t)=1$ for any $t$. Arrival rates are time-varying such that in the first $5000$ periods, there is a new type-$1$ post for each period ( $\lambda_{1}(t)=1$ and $\lambda_{2}(t)=0$ for $t\leq 5000$); and in the remaining periods, a type-$1$ post arrives with probability $0.6$ and a type-$2$ post arrive with probability $0.4$ ($\lambda_{1}(t)=0.6$ and $\lambda_{2}(t)=0.4$ for $t>5000$). We simulate OLBACID and BACID.UCB, the Optimism-Only approach introduced in Section 4.1, with $\beta=\gamma=1/\sqrt{K\ell_{\max}}$ as in Theorem 3. To set the confidence bounds in (11) and (12), we take $r_{\max}=1$ and $\sigma_{\max}=1$.
Proposition 3 shows that with high probability, BACID.UCB reviews zero type-$2$ posts because the queue of type-$1$ posts is always longer than that of type-$2$ posts. Our simulation confirms these observations as in Figures 6 and 6. In Figure 6, the curve for the number of reviewed type-$2$ posts under BACID.UCB disappears since it is zero for any period. The lack of reviewed samples translates to incorrect estimation for $r_{k}^{O}-r_{k}^{R}$, and thus wrong classification. In particular, Figure 7 shows that BACID.UCB fails to identify that $h_{2}$ is positive by maintaining an estimation of zero for $h_{2}$, while OLBACID is able to identify it early with label-driven admission and force scheduling.
Figure 5: Queue Lengths of both types
under BACID.UCB
Figure 6: Number of reviewed posts under BACID.UCB and OLBACID
Figure 7: Evolution of confidence bounds for $h_{k}=r_{k}^{O}-r_{k}^{R}$ under BACID.UCB and OLBACID
The error in classification leads to a larger regret of BACID.UCB compared to OLBACID as shown in Figure 9. However, such increase in regret depends on the difference of lifetimes between the two types of posts. On the one hand, when the lifetime of a type-$2$ post is really short, a mis-classification of this post will not lead to a large increase in regret. On the other hand, if the lifetime of a type-$2$ post is close to that of a type-$1$ post, BACID.UCB may not fail to review type-$2$ posts. Indeed, by varying the lifetime of type-$2$ posts as a fraction of that of type-$1$ posts (which is fixed to $10000$), our result in Figure 9 shows that the increase in regret behaves as an inverted-U function with respect to the ratio $\ell_{2}/\ell_{1}$. Specifically, the difference first increases linearly with respect to $\ell_{2}$ since in this region, BACID.UCB fails to classify every type-$2$ post and the cost of such mis-classification is linear in $\ell_{2}$. Nevertheless, when $\ell_{2}$ is large enough, BACID.UCB starts to admit and review type-$2$ posts and thus the difference in regret starts to decrease to 00.
Figure 8: $\textsc{Reg}^{\pi}(1,t)$ for periods $t\leq T$
[…]
Figure 9: Scaling in $\ell_{2}/\ell_{1}$ with $\ell_{1}=10^{4}$ for $\textsc{Reg}^{\textsc{BACID.UCB}}(1,T)-\textsc{Reg}^{\textsc{OLBACID}}(1,T)$